OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

Nnesha (nnesha):

set it equal to zero then you will get quadratic equation

OpenStudy (anonymous):

Yeah I noticed that but can you show me how you to set it up

Nnesha (nnesha):

\[\huge\rm 7\sin^2x-14sinx+2+5=0\] add 5 both

Nnesha (nnesha):

to make it easy let sin(x) x \[\huge\rm 7x^2 -14x+2+5 =0\]  solve for x :-)